### v1

# AppellateGen: A Benchmark for Appellate Legal Judgment Generation

Abstract
Legal judgment generation is a critical task in legal intelligence. However, existing research in legal judgment generation has predominantly focused on first-instance trials, relying on static fact-to-verdict mappings while neglecting the dialectical nature of appellate (second-instance) review. To address this, we introduce AppellateGen, a benchmark for second-instance legal judgment generation comprising 7,351 case pairs. The task requires models to draft legally binding judgments by reasoning over the initial verdict and evidentiary updates, thereby modeling the causal dependency between trial stages. We further propose a judicial Standard Operating Procedure (SOP)-based Legal Multi-Agent System (SLMAS) to simulate judicial workflows, which decomposes the generation process into discrete stages of issue identification, retrieval, and drafting. Experimental results indicate that while SLMAS improves logical consistency, the complexity of appellate reasoning remains a substantial challenge for current LLMs. The dataset and code are publicly available at: https://anonymous.4open.science/r/AppellateGen-5763.
AppellateGen: A Benchmark for Appellate Legal Judgment Generation
Hongkun Yang1,2thanks: Equal contribution. Work done while Hongkun Yang was an intern at Nanyang Technological University., Lionel Z. Wang1,3thanks: Equal contribution. Project Leader., Wei Fan4, Yiran Hu5, Lixu Wang1, Chenyu Liu3, Shenghong Fu3, Haoyang Li3, Xin Xu3, Jiexin Zheng6, Wei Dong1thanks: Corresponding author. 1Nanyang Technological University, 2Ocean University of China, 3The Hong Kong Polytechnic University, 4Hong Kong University of Science and Technology, 5Tsinghua University, 6Peking University
1 Introduction
The integration of Large Language Models (LLMs) into the legal domain has advanced legal NLP, moving from simple clause retrieval to complex reasoning tasks such as passing the Bar Exam (katz2024gpt), and transitioning from discriminative Legal Judgment Prediction (LJP) (cui2023chatlaw; feng2022legal) to Legal Document Generation (li2025casegen; yue2023disc). Despite these advancements, contemporary methodologies predominantly adopt a First-Instance Perspective (zhong2018legal; liu2023ml; xu2020distinguish), where models synthesize judgments directly from static factual descriptions. However, this linear approach neglects a critical phase: the Appellate (Second-Instance) Review (illustrated in Appendix A). Distinct from the fact-to-verdict mapping of a first trial, appellate proceedings are inherently dialectical, necessitating a comparative analysis of the original judgment against appellant grievances to rectify errors in fact-finding or legal application (wechsler1977appellate).
This distinction is particularly pronounced in Civil Law jurisdictions. Here, appellate courts operate under the principle of continuation of trial, allowing for the re-examination of facts and the admission of new evidence (apple1995primer). This procedural feature necessitates models capable of processing dynamic factual updates. According to judicial statistics from the Supreme People’s Court of China, the reversal rate of second-instance cases consistently hovered around 10% to 11% (exceeding 637,000 cases) from 2021 to 2023 SPC2021; SPC2022; SPC2023. This substantial volume of modified verdicts underscores the complexity inherent in appellate adjudication, which demands complex reasoning to identify discrepancies and produce a binding judgment that affirms, reverses, or remands the initial ruling.
Despite its importance, existing legal benchmarks exhibit limitations in Task Formulation and Contextual Granularity. First, regarding formulation, existing benchmarks primarily assess discriminative capabilities through multiple-choice questions or classification chalkidis-etal-2022-lexglue; fei-etal-2024-lawbench. While recent initiatives have integrated generative tasks li-etal-2025-legalagentbench; li2024lexeval, they predominantly focus on drafting documents from static facts, overlooking the dialectical nature of appeals. Second, in terms of granularity, current methodologies typically model legal documents as isolated data points. Retrieval-centric benchmarks pipitone2024legalbench and document generation datasets li2025casegen generally treat case processing as a single-step inference task. This approach neglects the procedural interdependency between adjudicatory stages, specifically the dependency between initial judgments and appellate reviews.
To bridge this gap, we introduce AppellateGen, the first benchmark dedicated to Second-Instance Legal Judgment Generation. Unlike previous benchmarks, AppellateGen models the complete dispute lifecycle. We curate 7,351 paired cases, linking first-instance judgments with their corresponding appellate outcomes. We enhance this with a fine-grained annotation schema that identifies points of contention, extracts new facts introduced during the appeal, and categorizes the rationale for reversal.
To validate the challenge posed by AppellateGen and investigate robust solutions, we propose the judicial Standard Operating Procedure (SOP)-based Legal Multi-Agent System (SLMAS) as a strong baseline. We observe that standard LLMs often struggle when processing the conflicting narratives and long-context evidence inherent in appellate review, leading to hallucinations. To address this, SLMAS imposes a Standard Operating Procedure (SOP) inspired by judicial workflows, decomposing the generation task into discrete stages: Analysis, Search, Predict, and Write. By orchestrating agents specialized in issue identification, retrieval, prediction, and drafting, our framework explicitly models the intermediate reasoning states required for valid adjudication. Empirical results demonstrate that this modular approach significantly enhances logical consistency and reversal prediction accuracy, surpassing both commercial and domain-specific baselines.
In summary, our contributions are threefold:
  • Pioneering Task Definition: We extend the frontier of legal NLP from static first-instance judgment prediction to Second-Instance Legal Judgment Generation. This task necessitates a paradigm shift towards dynamic dialectical reasoning, requiring models to identify judicial errors and synthesize conflicting narratives between the original verdict and new evidence.
  • The AppellateGen Benchmark: We construct the first large-scale benchmark of 7,351 paired legal cases that explicitly links trial proceedings with appellate outcomes. Equipped with a fine-grained annotation schema for reversal rationales and evidentiary updates, this dataset serves as a rigorous testbed for evaluating long-context reasoning and causal consistency in legal adjudication.
  • Methodological Framework & Insights: We propose a judicial Standard Operating Procedure (SOP)-based Legal Multi-Agent System (SLMAS). By decomposing the generation process into dispute identification, retrieval, prediction, and drafting, our framework significantly mitigates logical hallucinations. Furthermore, our extensive evaluation reveals that general-purpose LLMs with strong reasoning capabilities outperform traditional domain-specific models on this complex task, offering new directions for legal AI development.
2 Related Work
2.1 Legal Judgment Prediction and Case Document Generation
Legal Judgment Prediction (LJP) has progressed from utilizing hierarchical legal structures (luo2017learning) and few-shot learning (hu2018few) to deploying attention-based architectures (chen2019charge; li2019mann; kang2019creating) and LLMs (nigam2024rethinking) for outcome deduction. Despite these advancements, the field has predominantly prioritized prediction over full document generation. While recent initiatives like CaseGen (li2025casegen) have pivoted toward generation, they remain restricted to first-instance trials, neglecting the appellate stage. Within the Chinese judicial framework, second-instance judgments are pivotal for their supervisory role, correcting errors from lower courts and unifying adjudication standards. To address this gap, we introduce the task of Second-instance Legal Judgment Generation.
First Instance Second Instance
Full Text: A case concerning a dispute over the right of subrogation between the plaintiff, PICC, and the defendants, Mr. Wen… Full Text: Mr. Wen appealed the first-instance judgment in a case concerning a subrogation claim dispute…
Facts: Wen (fully liable) collided with Han. Han’s vehicle was repaired for \$8,586. Insurer paid Han and sought recovery from Wen. Wen claimed "old damages" existed but provided no proof. New Facts: Court retrieved police body-cam footage. Footage confirms Han admitted to "old injuries" on site. Damage height didn’t match the accident impact.
Legal Articles: - Insurance Law Art. 60 - Road Traffic Safety Law Art. 76 Legal Articles: - Civil Procedure Law Art. 170(1)(2)
Judgment: 1. PICC (Insurer 2) to pay \$2,000. 2. Wen (Defendant) to pay \$6,586. Judgment: 1. Affirm PICC’s \$2,000 payment. 2. Revoke Wen’s \$6,586 payment order.
Is_reversal: True  |  Reason for Reversal: Based on newly obtained police footage, the repair scope clearly exceeded the damage caused by the current accident. The trial court’s erroneous fact-finding regarding the damage amount is corrected, and the excess claim is dismissed.
Figure 1: An example from our dataset (translated from Chinese). Facts refers to the key factual findings extracted from the unstructured full text via a Large Language Model (LLM). Reason for Reversal denotes the explanatory rationale synthesized by the LLM, which identifies the critical discrepancies between the first and second instances that led to the judgment change.
2.2 Benchmarks in the Legal Domain
Existing benchmarks predominantly focus on isolated, single-stage tasks, varying significantly in scope. On an international scale, foundational resources like LegalBench (guha2023legalbench) and LexGLUE (chalkidis-etal-2022-lexglue) provide extensive classification annotations, while LEXTREME (niklaus2023lextreme), FairLex (chalkidis-etal-2022-fairlex), and ArabLegalEval (hijazi-etal-2024-arablegaleval) extend evaluation to multilingual and fairness contexts. In the domain of Chinese law, the CAIL competition series serves as a backbone for recent benchmarks emphasizing cognitive reasoning and syllogism, such as LawBench (fei-etal-2024-lawbench), LexEval (li2024lexeval), and LAiW (dai-etal-2025-laiw). Beyond general understanding, recent works have targeted specific system components, including retrieval (LegalBench-RAG (pipitone2024legalbench), LegalSearchLM (kim2025legalsearchlm)), instruction tuning (LawInstruct (niklaus-etal-2025-lawinstruct)), and agentic workflows (LegalAgentBench (li-etal-2025-legalagentbench)). While generative capabilities are partially addressed by CaseGen (li2025casegen) and GreekBarBench (chlapanis2025greekbarbench), these tasks often lack the dialectical depth of appellate review. A notable exception is the recent AppealCase dataset (huang2025appealcase), which explicitly targets second-instance scenarios. However, this work focuses on the discriminative classification task, failing to capture the complex reasoning required to draft appellate opinions.
2.3 Legal Multi-Agent Systems
Multi-Agent Systems (MAS) have addressed single-agent limitations such as hallucination through collaborative simulation (ji2023towards; li2023camel). In the legal domain, frameworks like ChatLaw (cui2023chatlaw) and logic-graph-enhanced models (yuan2026multi) leverage such collaboration to optimize first-instance judgment prediction and consultation. While these systems efficiently decompose tasks via role-playing, they are inherently designed for mapping facts to verdicts rather than the dialectical reasoning required for appellate review. Second-instance generation demands a comparative analysis to weigh original judgments against appeal grounds.
Figure 2: SOP-based Legal Multi-Agent System (SLMAS). With first-instance judgment and new facts provided, the framework mimics the human judicial Standard Operating Procedure (SOP) by orchestrating four specialized agents. These agents sequentially identify disputed issues, retrieve relevant legal statutes, predict the reversal outcome, and finally draft the appellate judgment, ensuring the generation is grounded in explicit legal reasoning.
3 AppellateGen Construction
3.1 Data Collection and Legal Instances Matching
We collected 4,706,987 raw legal judgment documents spanning from 2021 to 2024111This specific timeframe was selected to ensure legal consistency, as the Civil Code and the 11th Amendment to the Criminal Law of the People’s Republic of China were enacted in 2021, providing a stable statutory basis for the judgments in our dataset. from China Judgments Online [^2] https://wenshu.court.gov.cn/, the official government repository for Chinese judicial documents. Since the raw repository stores judgments as independent documents without explicit links between first-instance and second-instance proceedings, we developed a multi-stage pipeline to match the same cases in different instances (Details illustrated in Appendix B.1). To assess the quality of the matching process, we randomly sampled 5% of the paired cases for manual expert review, achieving a consistency rate of approximately 96%. After the whole process, all the documents have been de-identified to protect the privacy of litigants. We show an example of our dataset in Figure 1.
3.2 Dataset Statistics
We initially collected approximately 4 million raw legal documents through the pipeline described above. A final dataset comprising 7,351 high-quality case pairs is curated, this sparsity is attributed to three main factors: 1) Natural Filtering: A significant portion of cases conclude at the first instance without appeal; 2) Procedural Outcomes: In the appellate stage, a high volume of cases are resolved through mediation or withdrawal, which typically yield brief procedural rulings rather than the substantive judgments required for reasoning tasks; 3) Strict Quality Control: To ensure the reliability of the benchmark, we enforced rigorous matching criteria based on case numbers and litigant names, discarding pairs with incomplete chains of custody or insufficient textual content. We report the key descriptive statistics in Table 1.
Metric 1st Instance 2nd Instance
Count 7,351 7,351
Avg. Length 5,004.1 5,752.8
Max. Length 40,219 41,732
Min. Length 916 879
Label Dist. Rev.: 10.7%  Aff.: 89.3%
Table 1: Descriptive statistics of the dataset. Lengths are measured in Chinese characters. The dataset contains 7,351 pairs in total (10.7% Reversal, 89.3% Affirmation).
3.3 Fine-grained Data Annotation
While standard metrics (e.g., ROUGE (lin2004rouge)) assess lexical overlap, they fail to capture the legal validity and logical consistency required for the task of second-instance document generation. To address this limitation, we designed a fine-grained annotation schema for the matched case pairs. These annotations establish a rigorous semantic ground truth that enables multi-dimensional evaluation beyond surface-level text generation.
3.3.1 Annotation Schema
Unlike previous datasets that focus solely on static judgment outcomes, our schema captures the dynamic interaction and adjudicative evolution between the two instances. This allows us to assess whether the generated documents accurately reflect the legal reasoning shifts from the first to the second trial. The annotations were obtained via a hybrid framework combining LLM-driven automated extraction with strict human verification. The schema comprises four key dimensions:
Judgment Result Classification:
This binary label serves as the high-level outcome indicator. We classify the second-instance ruling as either Affirm (maintaining the original verdict) or Reverse (encompassing cases remanded for retrial or where the judgment was directly altered). This provides the fundamental stance for the generated text.
Rationale for Reversal:
For cases involving a reversal, we identify the specific legal trigger. This label distinguishes whether the appellate decision was driven by erroneous fact-finding, misapplication of law, or procedural violations in the first instance. This annotation is critical for evaluating if the generated document correctly explains the why behind the verdict.
New Evidence and Facts Identification:
This label summarizes newly ascertained facts derived exclusively from the second-instance judgment, explicitly excluding facts established in the first instance. This design requires the LLM to synthesize the first-instance factual summary with these novel appellate details to perform legal reasoning, thereby simulating authentic judicial practice.
Statutory Citation Extraction:
This label targets the specific legal articles cited by the appellate court as the controlling authority for its decision. We extract the full citation (e.g., Article X of the Civil Code), distinguishing operative laws that dictate the ruling from those merely referenced in the procedural history.
3.3.2 Automated Annotation Pipeline
To guarantee the reliability of annotations within complex legal contexts, we established a robust three-stage pipeline:
Structural Decomposition:
We employed customized regular expressions to partition unstructured judgment documents into distinct functional units, specifically the Factual Description and Judicial Reasoning sections. This segmentation mitigates input noise and directs the model’s attention toward contextually relevant information.
LLM-Driven Annotation:
We utilized DeepSeek-V3.2 liu2025deepseek to generate preliminary annotations. To bolster reasoning transparency and label accuracy, we incorporated a Chain-of-Thought (CoT) prompting strategy (Detailed in Appendix B.2). This approach compels the model to articulate the underlying legal rationale and calculate a self-assessed confidence score ($\gamma$) prior to determining the final label.
Human-in-the-Loop Verification:
To ensure the construction of a rigorous ground truth, we instituted a confidence-aware review protocol. Instances where the model’s confidence fell below a designated threshold ($\gamma<0.8$) were flagged for manual adjudication by legal experts.
4 SOP-based Legal Multi-Agent System
4.1 Problem Formulation
We formalize the task of Second-Instance Legal Judgment Generation as a conditional text generation problem involving multi-hop reasoning. Let $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ denote the dataset. For each case, the input context $\mathbf{x}$ comprises a tuple $\mathbf{x}=\langle\mathbf{d}_{\mathrm{1st}},\mathbf{e}_{\mathrm{new}}\rangle$, where $\mathbf{d}_{\mathrm{1st}}$ represents the textual content of the first-instance judgment (including initial fact-finding and verdict), and $\mathbf{e}_{\mathrm{new}}$ denotes the new evidentiary facts introduced during the appeal (where $\mathbf{e}_{\mathrm{new}}=\emptyset$ if no new evidence is presented).
The target output is the second-instance judgment document, denoted as $\mathbf{y}=(y_{1},y_{2},\dots,y_{T})$. Our objective is to learn a parameterized model $\theta$ that maximizes the log-likelihood of generating $\mathbf{y}$ conditioned on $\mathbf{x}$. Unlike standard end-to-end generation, we posit that the generation of $\mathbf{y}$ strictly depends on a sequence of intermediate reasoning states $\mathbf{z}=\{\mathbf{z}_{\mathrm{issue}},\mathbf{z}_{\mathrm{law}},\mathbf{z}_{\mathrm{verdict}}\}$. Thus, the objective function is formulated as:
$\mathcal{L}(\theta)=\sum_{(\mathbf{x},\mathbf{y})\in\mathcal{D}}\log P_{\theta}(\mathbf{y}\mid\mathbf{x},\mathbf{z})$ (1)
where $\mathbf{z}$ represents the latent rationales derived from the judicial Standard Operating Procedure (SOP).
4.2 Framework Overview
Directly mapping $\mathbf{x}\to\mathbf{y}$ often leads to hallucinations and logical inconsistencies due to the complexity of appellate review. To address this, we propose a judicial Standard Operating Procedure (SOP)-based Legal Multi-Agent System (SLMAS) (Shown in Figure 2) that mimics the human judicial Standard Operating Procedure (SOP) (demarco2013peopleware; hong2023metagpt).
We model the appellate process as a deterministic assembly line executed by four specialized agents:
$\mathcal{A}=\{\mathcal{A}_{\mathrm{issue}},\mathcal{A}_{\mathrm{retr}},\mathcal{A}_{\mathrm{pred}},\mathcal{A}_{\mathrm{write}}\}.$ (2)
Each agent $\mathcal{A}_{k}$ acts as a functional operator that transforms the current state and prior contexts into a structured intermediate output. The overall workflow forms a directed acyclic graph (DAG) of reasoning (digitale2022tutorial), ensuring that the final generation is grounded in explicit legal logic rather than statistical correlations.
4.3 Sequential Reasoning Workflow
The execution of our framework follows a sequential chain corresponding to the four stages of human adjudication: Analysis, Search, Predict, and Write.
Stage 1: Disputed Issues Identification (Analysis).
The process begins with the Disputed Issues Identification Agent ($\mathcal{A}_{\mathrm{issue}}$). This agent analyzes the discrepancy between the first-instance judgment and the new evidence to distill the core points of contention. Formally, given the input $\mathbf{x}$, the agent extracts a set of disputed issues $\mathbf{z}_{\mathrm{issue}}$:
$\mathbf{z}_{\mathrm{issue}}=\mathcal{A}_{\mathrm{issue}}(\mathbf{d}_{\mathrm{1st}},\mathbf{e}_{\mathrm{new}})$ (3)
where $\mathbf{z}_{\mathrm{issue}}$ serves as a structured summary of the appeal grounds, guiding the subsequent retrieval process.
Stage 2: Legal Retrieval (Search).
To ground the reasoning in statutory law, the Legal Retrieval Agent ($\mathcal{A}_{\mathrm{retr}}$) utilizes the identified issues $\mathbf{z}_{\mathrm{issue}}$ to query an external legal knowledge base $\mathcal{K}$ (Details Illustrated in Appendix C). Unlike generic retrieval, this agent focuses on statutes specifically relevant to the controversy:
$\mathbf{z}_{\mathrm{law}}=\mathcal{A}_{\mathrm{retr}}(\mathbf{z}_{\mathrm{issue}},\mathcal{K})$ (4)
Here, $\mathbf{z}_{\mathrm{law}}$ represents the retrieved relevant legal articles and judicial interpretations, providing the normative basis for the review.
Stage 3: Reversal Prediction (Predict).
Before drafting the document, the Reversal Prediction Agent ($\mathcal{A}_{\mathrm{pred}}$) mimics a judge’s internal deliberation. This agent integrates the case context and derived legal grounds to adjudicate the appeal, specifically predicting whether the primary verdict warrants affirmation or reversal. The agent generates a predictive rationale $\mathbf{z}_{\mathrm{verdict}}$:
$\mathbf{z}_{\mathrm{verdict}}=\mathcal{A}_{\mathrm{pred}}(\mathbf{d}_{\mathrm{1st}},\mathbf{e}_{\mathrm{new}},\mathbf{z}_{\mathrm{issue}},\mathbf{z}_{\mathrm{law}})$ (5)
This step acts as a logical checkpoint, ensuring the final document’s conclusion is consistent with the analyzed facts and laws.
Stage 4: Judgment Generation (Write).
Finally, the Document Writing Agent ($\mathcal{A}_{\mathrm{write}}$) generates the final second-instance judgment. Conditioned on the entire reasoning chain $\mathbf{z}=\{\mathbf{z}_{\mathrm{issue}},\mathbf{z}_{\mathrm{law}},\mathbf{z}_{\mathrm{verdict}}\}$, the agent produces the document token-by-token:
$P(\mathbf{y}\mid\mathbf{x},\mathbf{z})$ $=\prod_{t=1}^{T}P_{\theta}\bigl(y_{t}\mid y_{<t},\mathbf{d}_{\mathrm{1st}},\mathbf{e}_{\mathrm{new}},$ (6)
$\qquad\qquad\mathbf{z}_{\mathrm{issue}},\mathbf{z}_{\mathrm{law}},\mathbf{z}_{\mathrm{verdict}}\bigr)$
By explicitly conditioning on $\mathbf{z}$, the framework ensures that the generated text strictly adheres to the identified issues, retrieved laws, and predicted verdict, effectively mitigating hallucination.
Model Standard Metrics LLM-as-a-Judge Scores Reversal Prediction Accuracy
ROUGE-L BERTScore Verdict Consistency Fact Consistency Legal Application Logical Reasoning Average
Single Non- Reasoning Agent
Qwen3-235B-A22B-Instruct-2507 0.344 0.866 2.136 2.098 2.503 2.291 2.257 61.42%
Qwen3-30B-A3B-Instruct-2507 0.360 0.863 2.144 2.322 2.504 2.120 2.273 60.43%
Qwen3-8B 0.313 0.840 1.564 2.077 2.151 1.469 1.815 57.92%
Gemini-2.5-Flash-Lite 0.521 0.836 1.955 2.314 2.411 1.713 2.098 57.58%
Gemini-2.5-Flash 0.536 0.856 2.396 2.451 2.558 2.323 2.432 66.53%
Single Reasoning Agent
QwQ-32B 0.277 0.851 2.366 2.541 2.689 2.243 2.460 60.06%
Qwen3-30B-A3B-Thinking-2507 0.288 0.836 2.213 2.373 2.606 2.000 2.298 64.81%
DeepSeek-R1-0528-Qwen3-8B 0.316 0.837 2.137 2.229 2.293 1.697 2.089 57.87%
gpt-oss-120b 0.320 0.797 1.262 2.078 2.059 1.749 1.787 41.81%
Single Domain Agent
DISC-LawLLM 0.080 0.660 1.643 2.075 2.237 1.683 1.910 35.22%
Wisdom Interrogatory 0.333 0.721 1.443 1.983 1.731 1.240 1.599 37.34%
Multi Agents (Our SLMAS)
Qwen3-235B-A22B-Instruct-2507 0.337 0.878 2.353 2.472 2.705 2.615 2.536 67.24%
Qwen3-30B-A3B-Instruct-2507 0.353 0.834 2.460 2.406 2.430 2.274 2.393 66.49%
Qwen3-8B 0.272 0.810 1.712 2.181 2.308 1.684 1.971 61.27%
Table 2: Performance comparison on Appellate Legal Judgment Generation task. Bold indicates best performance.
5 Experiment
5.1 Experiment Settings
Test Dataset Construction.
We curated a dedicated test set through stratified sampling. We then obtained 1,000 distinct cases from the original dataset, consisting of 500 cases where the initial judgment was reversed and 500 cases where it was affirmed. This 1:1 ratio ensures that the evaluation metrics accurately reflect the model’s reasoning ability rather than the label distribution.
Baseline.
We leverage LLMs ranging from tens of billions to hundreds of billions of parameters as baselines. The models include: (1) General Open-source Models: gpt-oss-120b agarwal2025gpt and Qwen3 yang2025qwen3; (2) Commercial Close-source Models: Gemini 2.5 comanici2025gemini; (3) Legal-specific Models: Domain-adapted models including DISC-LawLLM yue2023disc and Wisdom Interrogatory wisdomInterrogatory.
Implementation Details.
All agents in SLMAS are instantiated using Qwen3 yang2025qwen3, with role-specific prompts and tool to enable functional specialization. For reproducibility, we fix the decoding temperature at 0.1 across all trials. To ensure fair comparison, identical prompts are employed, and input sequences are truncated to fit within each model’s respective context window.
5.2 Automatic Evaluation
We employ a hybrid framework combining LLM-as-a-Judge (reference-free expert scoring) li2025generation and standard reference-based metrics to assess the legal validity of generated judgments.
LLM-as-a-Judge.
We utilize DeepSeek-V3.2 liu2025deepseek to score generations on a Likert scale (0–5) based on expert-designed rubrics joshi2015likert. The detailed rubrics are located in Appendix D. The evaluation encompasses four dimensions: (1) Verdict Consistency: Checks if the ruling direction (Affirm/Reverse) matches the ground truth and whether the judgment results are the same. A mismatch results in a nullified score ($S_{verdict}=0$). (2) Fact Consistency: Assesses the reconstruction of facts and handling of evidence. (3) Legal Application: Verifies the accurate citation of operative statutes. (4) Logical Reasoning: Evaluates syllogistic rigor regarding appeal grounds.
Standard Metrics.
Complementing the judge model, we report ROUGE-L lin2004rouge and BERTScore zhang2019bertscore to measure lexical overlap and semantic similarity against reference judgments.
Agent Configuration Standard Metrics LLM-as-a-Judge Scores Reversal Prediction Accuracy
Disp. Issues Legal Retr. Rev. Pred. Doc. Gen. ROUGE-L BERTScore Verd. Cons. Fact Cons. Legal App. Logical Reas. Average
0.272 0.810 1.712 2.181 2.308 1.684 1.971 61.27%
0.278 0.820 1.694 2.064 2.085 1.644 1.872 59.30%
0.269 0.803 1.663 2.129 2.191 1.651 1.909 60.41%
0.331 0.823 1.621 2.118 2.131 1.614 1.871 58.05%
0.313 0.840 1.564 2.077 2.151 1.469 1.815 57.92%
Table 3: We perform the Ablation study using Qwen3-8B. We report Generation Similarity (ROUGE-L) and (BERTScore), LLM-as-a-Judge Scores (including Verdict Consistency, Fact Consistency, Legal Application and Logical Reasoning), and Reversal Prediction Accuracy. The Average column represents the mean of the LLM-as-a-Judge Scores. Best results are bold.
5.3 Main Result
5.3.1 Effectiveness of the SLMAS Framework
Table 2 demonstrates that the SLMAS framework consistently enhances performance across models of varying scales. Using Qwen3-235B-A22B-Instruct-2507 as a primary example, SLMAS improves the average generation quality score from 2.257 to 2.536 and increases Reversal Prediction Accuracy from 61.42% to 67.24%. This performance exceeds that of the proprietary Gemini 2.5 Flash (66.53%), indicating that open-source models can achieve state-of-the-art results when equipped with effective agentic strategies. The improvement is attributed to the structural decomposition of the legal syllogism. By decoupling issue identification from the drafting process, SLMAS effectively models the judicial workflow, thereby mitigating the hallucinations and logical inconsistencies observed in long-context generation tasks.
5.3.2 Reasoning Capabilities versus Domain Knowledge
A significant divergence in performance is observed between general-purpose LLMs and domain-specific models. While general models achieve accuracies exceeding 60%, traditional legal models such as DISC-LawLLM and Wisdom Interrogatory lag significantly, with accuracies between 35% and 37%. This disparity suggests that appellate judgment generation relies less on static knowledge retrieval and more on comparative reasoning capabilities. Domain-specific models, trained on static legal corpora, lack the cognitive flexibility required to process dynamic evidentiary conflicts between first and second-instance trials. In contrast, general LLMs, particularly those enhanced with chain-of-thought processes, demonstrate superior adaptability in these logic-intensive scenarios.
5.3.3 Persistent Challenges in Generation Quality
Despite the improvements introduced by SLMAS, generating second-instance judgments remains a complex challenge. Although prediction accuracy has reached approximately 67%, the generation quality scores reveal a persistent gap, with average scores hovering around 2.5 out of 5. The primary bottleneck is identified in the Logical Reasoning metric (2.615), where models occasionally fail to construct closed-loop arguments that rigorously connect new evidence to statutory conclusions. This indicates that while instruction tuning improves adherence to format, future advancements may require paradigms specifically designed to enhance long-chain causal reasoning.
5.4 Ablation Study
To evaluate the contribution of individual components, we conducted an ablation study using Qwen3-8B on the test dataset. The results, presented in Table 3, highlight the interdependency of reasoning modules and the divergence between surface-level metrics and semantic validity.
Dependency of Retrieval on Issue Identification
The full system achieves optimal performance with an average score of 1.971 and a reversal prediction accuracy of 61.27%. A critical finding is the impact of the Disputed Issues Identification agent. Removing this module degrades performance not only in reasoning but also in legal application. Without explicit issue definition, the subsequent retrieval module lacks directional focus, leading to the retrieval of statutes that fail to ground the final judgment. This confirms that SLMAS prevents error propagation by ensuring that each stage is conditioned on a verified intermediate state.
Divergence between N-gram and Semantic Metrics
An analysis of the No Agents baseline reveals a discrepancy between traditional metrics and judge-based evaluations. This configuration yields the highest ROUGE-L score (0.331) but a mediocre LLM-as-a-Judge score (1.871). This anomaly exposes the limitation of surface-level metrics in legal generation tasks. The baseline model tends toward verbatim repetition of the first-instance judgment, resulting in high lexical overlap but low dialectical value. Conversely, SLMAS promotes logical restructuring to address appeal grounds, which naturally reduces n-gram overlap with the source text while significantly improving Verdict Consistency (1.712 vs. 1.621) and Logical Reasoning. This underscores the necessity of employing semantic-aware evaluation over traditional overlap metrics for high-level reasoning tasks.
6 Conclusion
We present AppellateGen, a benchmark shifting the focus of legal intelligence from judgment prediction to the dialectical complexity of second-instance judgment generation. Comprising 7,351 annotated case pairs, it evaluates model reasoning over conflicting narratives and judicial errors. Our proposed SLMAS framework, which structurally decomposes the judicial workflow, significantly enhances logical consistency and reversal accuracy. Notably, general LLMs equipped with reasoning capabilities outperform domain-specific models, underscoring the primacy of cognitive flexibility over static domain pre-training for this task. While a gap remains between generated outputs and human expertise, AppellateGen establishes a foundation for future research into causal dependency modeling and robust comparative reasoning architectures.
Limitations and Future Work
Our study presents a novel benchmark for second-instance legal judgment generation. However, we acknowledge two primary limitations. First, the geographic and linguistic scope of AppellateGen is confined to Mainland China and the Chinese language. This jurisdiction-specific focus, while necessary for legal precision, limits the direct applicability of our models to other legal systems or multilingual contexts. Second, a significant performance bottleneck persists. Although our proposed SLMAS outperforms most of the baselines, the absolute performance remains suboptimal, with a SOTA metric of $67.24\%$ and a LLM-as-a-judge evaluation score of only $2.536$ out of $5$. These results suggest that current paradigms struggle to fully capture the intricate dialectical reasoning required for high-stakes judicial writing. To address these challenges, we look forward to the emergence of foundation models possessing superior legal reasoning capabilities.
Ethics Statement
The judgment documents of our dataset are all collected from a publicly accessible platform, China Judgments Online, which is widely used in legal AI research huang2025appealcase; li2025casegen. We strictly adhere to the data usage policies and redistribution terms of the source platform, ensuring that our release is confined to de-identified textual content within the boundaries of public access. However, legal data contains sensitive information. To strictly protect the privacy of the individuals involved, we have applied rigorous anonymization preprocessing. All Personally Identifiable Information (PII), such as litigants’ names, ID numbers, home addresses, and phone numbers, has been desensitized prior to model training and dataset release. We also acknowledge the potential risks associated with algorithmic bias. As the dataset reflects historical judicial decisions, it may inherently contain biases related to gender, region, or specific crime types. Models trained on this data might perpetuate these biases. Therefore, we emphasize that this dataset and the proposed method are intended for academic research and as assistive tools for legal professionals only. They should never be used as a substitute for human judges or to automate final judicial decisions. We strongly advise future researchers to conduct fairness evaluations when deploying systems based on this dataset.
Appendix A Background on Appellate (Second-Instance) Review in Civil Law Jurisdictions
Figure 3: Proceedings of the Second Instance: Following the filing of an appeal by the appellant, the admission of new evidence may result in either the reversal of the verdict or the affirmation of the original judgment.
The procedural scope of appellate review exhibits fundamental differences between Common Law and Civil Law jurisdictions. In Common Law systems, appellate courts typically exercise deference to the lower court regarding factual determinations, restricting their review primarily to errors of law. Conversely, Civil Law jurisdictions adopt a comprehensive approach where the court of second instance functions as a continuation of the trial process, retaining the authority to re-examine both legal applications and factual findings. This study focuses specifically on the Civil Law framework, utilizing a dataset derived from Chinese legal proceedings.
Within this context, the second-instance trial serves as a critical remedial mechanism. As illustrated in Figure 3, the procedure is initiated when a litigant, who may be either the plaintiff or the defendant, files an appeal expressing dissatisfaction with the first-instance judgment. A defining characteristic of this process in China is the admissibility of new evidence. Unlike systems that strictly limit the appellate record, this framework permits the introduction of supplementary facts that were previously unavailable or overlooked.
Consequently, the input for the appellate judge comprises two distinct streams of information: the text of the original first-instance judgment and the set of newly presented evidentiary facts. The presiding judge is tasked with synthesizing these inputs to evaluate the validity of the appeal. This review process necessitates a comparative analysis where the judge assesses whether the new evidence sufficiently contradicts the original findings or if the initial application of the law remains robust. The proceeding culminates in a binary judicial determination. The court may decide to affirm the original judgment if the appeal lacks sufficient merit, effectively affirming the prior ruling. Alternatively, if the combination of the original context and new facts reveals substantive errors, the court issues a reversal, thereby altering the legal verdict. Figure 3 provides a schematic representation of this workflow, delineating the causal path from the filing of the appeal and the admission of evidence to the final adjudication.
Appendix B AppellateGen Construction Details
To construct AppellateGen, we perform two main steps: (1) Data Collection and Legal Instance Matching, and (2) Fine-grained Data Annotation. The construction details are as follows.
B.1 Legal Instances Matching
First, we employed high-precision Regular Expressions (RE) to extract unique case identifier of first instance from second-instance judgment document. The specific extraction pattern is detailed in Figure 4. Subsequently, we retrieved the corresponding first-instance judgment documents based on these identifiers to construct candidate case pairs.
Figure 4: RE pattern for legal instances matching. This pattern can extract the case identifier of first instance from second-instance judgment document.
To ensure the semantic integrity of these pairs, filtering out potential mismatches caused by clerical errors in case numbers, we utilized a LLM as a semantic validator. Specifically, we prompted the LLM to compare the litigants, the cause of action, and the court hierarchy described in both documents, outputting a binary verification score indicating whether they refer to the identical legal dispute. The used prompts are detailed below.
You are an experienced expert in reviewing Chinese legal documents, specializing in the filing and correlation of civil, criminal, and administrative case files. You possess exceptional logical analysis and attention to detail. Your task is to determine whether two given legal documents (a first-instance judgment and a second-instance judgment) belong to the same case’s proceedings (i.e., whether the second instance appealed against the first-instance judgment). Please follow these steps for logical reasoning, without directly providing a conclusion: 1. Extract Key Information: From the first-instance judgment: Extract the first-instance case number, plaintiff/defendant’s name (or title), and cause of action. From the second-instance judgment: Extract the second-instance case number, appellant/respondent’s name, and the original trial case number cited in the text. 2. Compare Relevance: Case Number Verification: Check the "Case Origin" or "Original Trial Circumstances" section… Verification of Parties: Does the appellant/appellee in the second instance correspond to the plaintiff/defendant… Verification of Facts: Are the core facts of the dispute consistent? 3. Elimination of Interference: Pay attention to distinguishing between "related cases"… Output Format: Please strictly follow the following JSON format for outputting results: { "is_same_case": true/false, "confidence_score": "0-10", "reasoning": "" }
B.2 Fine-grained Data Annotation
Here is the prompt template for LLM-driven annotation.
Please act as a professional expert in Chinese legal document analysis. I will provide the full text of the [First Instance Judgment] and the [Second Instance Judgment] for the same case. Your tasks are: 1. Deep Analysis and Extraction: Compare the two documents and extract key information. 2. Formatted Output: Integrate all analysis results into a strict JSON object. Please output strictly according to the following JSON structure and do not include Markdown tags: {
"reasoning_trace": "Your step-by-step analysis here...",
  "confidence_score": float,
  "is_reversal": bool,
  "first_instance": {
    "facts": "string",
    "disputed_issues": "string",
    "legal_articles": ["string"],
    "judgment": "string"
  "second_instance": {
    "new_facts": "string",
    "disputed_issues": "string",
    "legal_articles": ["string"],
    "judgment": "string"
  "reason_for_reversal": "string"
Here are the explanations of the following JSON schemas.
is_reversal: A boolean value indicating whether the second-instance judgment reverses. first_instance: A structured summary of the first-instance judgment, including factual findings, disputed issues, cited legal articles, and the final ruling. facts: Core facts established in the first instance. disputed_issues: Key disputes addressed by the court. legal_articles: Statutory provisions relied upon. judgment: Outcome of the first-instance ruling. second_instance: A structured summary of the appellate judgment, emphasizing changes relative to the first instance. new_facts: Newly identified facts or evidence. disputed_issues: Disputed issues considered on appeal. legal_articles: Controlling statutory provisions. judgment: Outcome of the second-instance ruling. reason_for_reversal: Legal grounds for reversal; left empty if the judgment is affirmed. Please begin analyzing the document content I provide.
Appendix C SLMAS details
We established a specialized legal knowledge base derived from statutes promulgated after 2021, utilizing the open-source LawRefBook toolkit LawRefBook2025. A key advantage of this repository is its capability to maintain synchronization with the authoritative National Laws and Regulations Database of China [^3] This platform serves as the official legislative database maintained by the Ministry of Justice of the People’s Republic of China (https://flk.npc.gov.cn).. To enable efficient semantic matching, the textual content of every legal article was embedded using Qwen3-Embedding zhang2025qwen3.
Upon receiving a search query, the retriever performs a coarse-grained search to identify the top-10 most similar candidates. To ensure precision, an LLM then functions as a fine-grained selector, filtering the retrieved results to pinpoint the specific legal articles and judicial interpretations applicable to the dispute.
Appendix D Evaluation Details
In this section, we provide the specific scoring rubrics and the prompt template used for the LLM-as-a-Judge evaluation described in Section 5.2.
Scoring Rubrics
We designed fine-grained criteria to guide the evaluator (DeepSeek-V3.2) in scoring the generated judgments. The scoring scale ranges from 0 to 5 across four specific dimensions. The detailed rubrics are presented in Table 4.
Evaluation Prompt Template
We employed a zero-shot prompting strategy with detailed instructions to ensure the LLM evaluator acts as an impartial judge. The full prompt used in our experiments is presented below.
System Role: You are a senior legal expert in the Chinese judicial system, specializing in appellate review. Your task is to evaluate a second-instance judgment. Evaluation Task: Please score the generated output from 0 to 5 on the following dimensions based on the provided rubrics. 1. Verdict Consistency (0-5): Does the model make the correct decision (Affirm vs. Reverse)? CRITICAL RULE: If the ruling direction contradicts the Ground Truth (e.g., model affirms while truth reverses), you MUST give 0 points. 2. Fact Consistency (0-5): Does the model accurately describe the facts? Crucially, does it capture the "New Evidence" introduced in the second instance? If the model ignores new evidence that leads to a reversal, score low. 3. Legal Application (0-5): Are the cited statutes correct and relevant? Does it cite the operative substantive law rather than just procedural descriptions? 4. Logical Reasoning (0-5): Is the reasoning coherent? Does it correctly refute or support the appeal grounds? For reversal cases, does it clearly explain why the first instance was wrong? Output Format: Please output a strictly valid JSON object: { "verdict_score": int, "fact_score": int, "law_score": int, "logic_score": int, "justification": "A brief explanation of the scoring…" }
Dimension Score Criteria Description
1. Verdict Consistency 0 Incorrect Ruling Direction. The generated judgment contradicts the ground truth (e.g., affirming instead of reversing). *Strict Penalty*
1–2 Ambiguous/Incomplete. The direction is unclear, contradictory, or missing the operative part entirely.
3–4 Direction Correct but Flawed. The general ruling (Affirm/Reverse) is correct, but specific execution orders (e.g., amounts, specific behaviors) are imprecise or missing.
5 Perfect Match. The operative part is identical in legal effect and scope to the ground truth, fully resolving the dispute.
2. Fact Consistency 0–1 Hallucination. Fabricates facts not present in the input or contradicts the core timeline of the case.
2–3 Missing New Evidence. Accurately recaps the first-instance facts but fails to identify or incorporate the "New Facts" introduced during the appeal.
4 Accurate. Captures both original facts and new evidence, with minor omissions in non-critical details.
5 Comprehensive. Perfectly reconstructs the factual basis, highlighting the contradiction between new evidence and original findings.
3. Legal Application 0–1 Erroneous Citation. Cites non-existent statutes or applies laws completely irrelevant to the case cause.
2–3 Incomplete Citation. Misses the key operative article that dictates the reversal/affirmation, citing only procedural rules.
4–5 Precise Citation. Correctly identifies and applies the specific articles (e.g., Civil Code Art. X) used in the ground truth.
4. Logical Reasoning 0–1 Incoherent. The reasoning is logically fractured or fails to address the appellant’s grievances entirely.
2–3 Mechanical Mapping. Performs simple fact-to-law mapping but lacks comparative analysis. Misses the causal link between new evidence and the verdict change.
4 Strong Syllogism. Explicitly discusses why the first-instance judgment was erroneous based on the points of contention.
5 Expert Dialectics. Demonstrates a clear "review-and-correct" logic, forming a watertight argument that mirrors a human judge’s opinion.
Table 4: Detailed Scoring Rubrics for Automatic Evaluation. The evaluator scores each generated judgment from 0 to 5 based on these criteria. Note that a score of 0 in Verdict Consistency is strictly enforced if the ruling direction is wrong.