### v2

# AppellateGen: A Benchmark for Appellate Legal Judgment Generation

Abstract
Legal judgment generation is a critical task in legal intelligence. However, existing research in legal judgment generation has predominantly focused on first-instance trials, relying on static fact-to-verdict mappings while neglecting the dialectical nature of appellate (second-instance) review. To address this, we introduce AppellateGen, a benchmark for second-instance legal judgment generation comprising 7,351 case pairs. The task requires models to draft legally binding judgments by reasoning over the initial verdict and evidentiary updates, thereby modeling the causal dependency between trial stages. We further propose a judicial Standard Operating Procedure (SOP)-based Legal Multi-Agent System (SLMAS) to simulate judicial workflows, which decomposes the generation process into discrete stages of issue identification, retrieval, and drafting. Experimental results indicate that while SLMAS improves logical consistency, the complexity of appellate reasoning remains a substantial challenge for current LLMs. The dataset and code are publicly available at: https://anonymous.4open.science/r/AppellateGen-5763.
AppellateGen: A Benchmark for Appellate Legal Judgment Generation
Hongkun Yang1,2thanks: Equal contribution.thanks: Work done while Hongkun Yang was an intern at Nanyang Technological University., Lionel Z. Wang1,311footnotemark: 1thanks: Project Leader., Wei Fan4, Yiran Hu5, Lixu Wang1, Chenyu Liu3, Shenghong Fu3, Haoyang Li3, Xin Xu3, Jiexin Zheng6, Wei Dong1thanks: Corresponding author. 1Nanyang Technological University, 2Ocean University of China, 3The Hong Kong Polytechnic University, 4Hong Kong University of Science and Technology, 5Tsinghua University, 6Peking University
1 Introduction
The integration of Large Language Models (LLMs) into the legal domain has advanced legal NLP, moving from simple clause retrieval to complex reasoning tasks such as passing the Bar Exam (katz2024gpt), and transitioning from discriminative Legal Judgment Prediction (LJP) (cui2023chatlaw; feng2022legal) to Legal Document Generation (li2025casegen; yue2023disc). Despite these advancements, contemporary methodologies predominantly adopt a First-Instance Perspective (zhong2018legal; liu2023ml; xu2020distinguish), where generative models synthesize judgments directly from static factual descriptions. However, this linear approach neglects a critical phase: the Appellate (Second-Instance) Review (illustrated in Appendix LABEL:app:_appellate). Distinct from the fact-to-verdict mapping of a first trial, appellate proceedings are inherently dialectical, necessitating a comparative analysis of the original judgment against appellant grievances to rectify errors in fact-finding or legal application (wechsler1977appellate).
This distinction is particularly pronounced in Civil Law jurisdictions. Here, appellate courts operate under the principle of continuation of trial, allowing for the re-examination of facts and the admission of new evidence (apple1995primer). This procedural feature necessitates models capable of processing dynamic factual updates. According to judicial statistics from the Supreme People’s Court of China, the reversal rate of second-instance cases consistently hovered around 10% to 11% (exceeding 637,000 cases) from 2021 to 2023 SPC2021; SPC2022; SPC2023. This substantial volume of modified verdicts underscores the complexity inherent in appellate adjudication (luneburg1981specially; bruhl2010deciding), which demands complex reasoning to identify discrepancies and produce a binding judgment that affirms, reverses, or remands the initial ruling.
Despite the importance of appellate review, existing legal benchmarks exhibit limitations in Task Formulation and Contextual Granularity. First, regarding formulation, existing benchmarks primarily assess discriminative capabilities through multiple-choice questions or classification chalkidis-etal-2022-lexglue; fei-etal-2024-lawbench. While recent initiatives have integrated generative tasks li-etal-2025-legalagentbench; li2024lexeval, they predominantly focus on drafting documents from static facts, overlooking the dialectical nature of appeals. Second, in terms of granularity, current methodologies typically model legal documents as isolated data points. Retrieval-centric benchmarks pipitone2024legalbench and document generation datasets li2025casegen generally treat case processing as a single-step inference task. This approach neglects the procedural interdependency between adjudicatory stages, specifically the dependency between initial judgments and appellate reviews.
To bridge this gap, we introduce AppellateGen, the first benchmark dedicated to Second-Instance Legal Judgment Generation. Unlike previous benchmarks, AppellateGen models the complete dispute lifecycle. We curate 7,351 paired cases, linking first-instance judgments with their corresponding appellate outcomes. We enhance this with a fine-grained annotation schema that identifies points of contention, extracts new facts introduced during the appeal, and categorizes the rationale for reversal.
To validate the challenge posed by AppellateGen and investigate robust solutions, we propose the judicial Standard Operating Procedure (SOP)-based Legal Multi-Agent System (SLMAS) as a strong baseline. We observe that standard LLMs often struggle when processing the conflicting narratives and long-context evidence inherent in appellate review, leading to hallucinations. To address this, SLMAS imposes a Standard Operating Procedure (SOP) inspired by judicial workflows, decomposing the generation task into discrete stages: Analysis, Search, Predict, and Write. By orchestrating agents specialized in issue identification, retrieval, prediction, and drafting, our framework explicitly models the intermediate reasoning states required for valid adjudication. Empirical results demonstrate that this modular approach significantly enhances logical consistency and reversal prediction accuracy, surpassing both commercial and domain-specific baselines.
In summary, our contributions are threefold:
  • Pioneering Task Definition: We extend the frontier of legal NLP from static first-instance judgment prediction to Second-Instance Legal Judgment Generation. This task necessitates a paradigm shift towards dynamic dialectical reasoning, requiring generative models to identify judicial errors and synthesize conflicting narratives between the original verdict and new evidence.
  • The AppellateGen Benchmark: We construct the first large-scale benchmark of 7,351 paired legal cases that explicitly links trial proceedings with appellate outcomes. Equipped with a fine-grained annotation schema for reversal rationales and evidentiary updates, this dataset serves as a rigorous testbed for evaluating long-context reasoning and causal consistency in legal adjudication.
  • Methodological Framework & Insights: We propose a judicial Standard Operating Procedure (SOP)-based Legal Multi-Agent System (SLMAS). By decomposing the generation process into dispute identification, retrieval, prediction, and drafting, SLMAS has more outstanding generation performance and mitigates logical hallucinations. Furthermore, our extensive evaluation reveals that general-purpose LLMs with strong reasoning capabilities outperform traditional legal domain-specific models on this complex task.
2 Related Work
2.1 Legal Judgment Prediction and Case Document Generation
Legal Judgment Prediction (LJP) has progressed from utilizing hierarchical legal structures (luo2017learning) and few-shot learning (hu2018few) to deploying attention-based architectures (chen2019charge; li2019mann; kang2019creating) and LLMs (nigam2024rethinking) for outcome deduction. Despite these advancements, the field has predominantly prioritized prediction over full document generation. While recent initiatives like CaseGen (li2025casegen) have pivoted toward generation, they remain restricted to first-instance trials, neglecting the appellate stage. Within the Chinese judicial framework, second-instance judgments are pivotal for their supervisory role, correcting errors from lower courts and unifying adjudication standards. To address this gap, we introduce the task of Second-instance Legal Judgment Generation.